The first difference of a sequence is 3, 6, 9, 12, 15...
Find the first six terms of the original sequence in each of the following cases:
a. The first term of the original sequence is 3.
b. The sum of the first two terms of the original sequence is 7.

Answer by chessace(471)

a. Just keep adding: 3 +3=6 +6=12 +9=21 ...
b. n is first number, n + n+3 = 7. 2n=4. n=2 and proceed similar to case a.